OpenStudy (anonymous):

Can you solve 5x - 6 = 3x - 8

OpenStudy (igreen):

5x - 6 = 3x - 8 FIrst, subtract 3x to both sides.

OpenStudy (igreen):

First*

OpenStudy (anonymous):

subtract 3x from both sides

OpenStudy (alexandervonhumboldt2):

first subtartc 3x from obth sides you get 5x-6-3x=3x-3x-8

OpenStudy (alexandervonhumboldt2):

you get 2x-6=-8

OpenStudy (anonymous):

x=-1

OpenStudy (alexandervonhumboldt2):

now add 6 to both sides

OpenStudy (alexandervonhumboldt2):

no direct ansers @Dubba2000

OpenStudy (alexandervonhumboldt2):

2x=-2 now just divide both sides by 2 and get your anseer

OpenStudy (anonymous):

@iGreen can u help me with my question

OpenStudy (anonymous):

i think @AlexandervonHumboldt2 deserves the medal he/she have a better explantion.

OpenStudy (alexandervonhumboldt2):

thx i'm a he

OpenStudy (anonymous):

can someone help me with my question

OpenStudy (alexandervonhumboldt2):

Tag